# Non-crossing permutations for the KP solitons under the Gel’fand-Dickey reductions and the vertex operators

### v1

@@ 1. Introduction
The KP equation is integrable, and has infinitely many symmetries represented by the commuting flows. The parameters of the flows are expressed by $\{t_{j}:j=1,2,\ldots\}$, in particular, we denote $t_{1}=x,t_{2}=y,t_{3}=t$ for the KP equation. The set of all the flows defines the KP hierarchy, whose first member is the KP equation, and the variable $u$ is now considered as a function of $\mathbf{t}=(t_{1},t_{2},\ldots)$. The real regular soliton solutions of the KP equation, referred to as the KP solitons, has been recently classified in [2, 9] (see the book [7] for the review of these results). Each KP soliton can be determined by a pair of two real data $(\kappa,A)$, referred to as soliton parameters, where $\kappa=(\kappa_{1},\ldots,\kappa_{M})\in\mathbb{R}^{M}$ and an $N\times M$ matrix $A\in\text{Gr}(N,M)_{\geq 0}$, the totally nonnegative (tnn) Grassmannain defined by the set of $N$-dimensional subspaces in $\mathbb{R}^{M}$ with all nonnegative maximal minors of $A$. We also recall that the element $A\in\text{Gr}(N,M)_{\geq 0}$ can be uniquely parameterized by the derangement (permutation without fixed point, i.e. no 1-cycle) of symmetric group $S_{M}$, denoted by $\pi(A)\in S_{M}$. Then the soliton solution $u(x,y,t)$ is given in terms of the $\tau$-function (see e.g. [7]),
@@
where 2-cycle $(i,j)$ implies the transposition $i\leftrightarrow j$. Note here that all the 2-cycles in (1.4) form a nesting (see below, also [3] for the definition). Each soliton with the parameter $\kappa=(\kappa_{k},-\kappa_{k})$ in the $N$-soliton solution is represented by the transposition $(k,2N-k+1)$, which is considered as the derangement associated to $A_{k}\in\text{Gr}(1,2)_{\geq 0}$, i.e. $\pi(A_{k})=(k,2N-k+1)$ and $\pi(A)=\prod_{k}\pi(A_{k})$.
@@
which is the spectral curve of solitons for the 3-reduction. As one of the main results (Theorem 4.2), we give the regular soliton solutions of the Boussinesq equation in (1.7), whose the soliton matrix $A$ is parametrized by the derangement in the cycle notation either
@@
where each $\kappa_{j}[\alpha_{i}]$ is a root of $\Phi_{\ell}(\kappa,\alpha_{i})=0$, and $\mathcal{I}[\alpha_{i}]$ is a subset of []:=$[\ell]:=\{1,\ldots,\ell\}$ with $|\mathcal{I}[\alpha_{i}]|=m_{i}$. Here $M=m_{1}+\cdots+m_{K}$. For each $m_{i}$, take $11\leq n_{i}\leq m_{i}-$ 1, and set $N=n_{1}+\cdots+n_{K}$. Then for each pair $(n_{i},m_{i})$, construct the matrix element $A[\alpha_{i}]\in\text{Gr}(n_{i},m_{i})_{\geq 0}$, so that the corresponding derangements $\pi(A[\alpha_{i}])$ are mutually non-crossing for $i=1,\ldots,K$. Then the regular soliton solution of the $\ell$-reduction is given by the soliton parameters $\kappa\in\mathbb{R}^{M}$ and $A[\alpha_{1},\ldots,\alpha_{K}]\in\text{Gr}(N,M)_{\geq 0}$ associated with a ($\kappa$-)direct sum of the matrices $A[\alpha_{i}]$,
^${\widehat{\bigoplus}}_{i=1}^{K}A[\alpha_{i}],$
@@ 2.2. $N$ truncation and the $\tau$-function
where $p_{i}({\bf x})$ is the elementary Schur polynomial of degree $i$ and ~$\tilde{\partial}=(\partial_{1},\frac{1}{2}\partial_{2},\frac{1}{3}\partial_{3},\ldots)$, which is defined by
@@ 2.3. KP solitons
where $\kappa\in\mathbb{R}^{M}$ whose elements are ordered as $(\kappa_{1}<\cdots<\kappa_{M})$, and $A=(a_{i,j})$ is an $N\times M$ real matrix with full rank. Let $I=\{i_{1}<i_{2}<\cdots<i_{N}\}$ be an $N$ ordered subset of the index set $[M]:=\{1,2,\ldots,M\}$. Then the $\tau$-function (2.6) becomes (1.3), and using the Binet-Cauchy lemma, the $\tau$-function can be expressed by
@@
    For $y\gg 0$, there are $[i_{n},\pi(i_{n})]$-solitons with $\pi(i_{n})>i_{n}$ for $n=1,\ldots,N$.
@@
    For $y\ll 0$, there are $[\pi(j_{m}),j_{m}]$-solitons with $\pi(j_{m})<j_{m}$ for $m=1,\ldots,M-N$.
@@ 2.4. The Gel’fand-Dickey $\ell$-reductions
[…]
@@
One should note that the reduction equation (2.12) implies that we have $f_{i}(\mathbf{t})=e^{\mathop{\textstyle\sum}\limits_{n}\alpha_{k}^{n}t_{n\ell}}g(\mathbf{t}^{\prime})$ with $\mathbf{t}^{\prime}=(t_{j}:j\neq 0~{}(\text{mod}~{}\ell))$, which shows that the $\tau$-function is given by $\tau(\mathbf{t})=\text{Wr}(f_{1},\ldots f_{N})(\mathbf{t})=e^{\mathop{\textstyle\sum}\limits_{k}\mathop{\textstyle\sum}\limits_{n}\alpha_{k}^{n}t_{n\ell}}\text{Wr}(g_{1},\ldots,g_{N})(\mathbf{t}^{\prime})$. Then the variables $r_{n}(\mathbf{t})$ in (2.11) has no dependency on the flow parameters $t_{n\ell}$.
@@ 3.1. The spectral curve and the $\tau$-function
(3.2) $t_{n}~{}\rightarrow~{}\left\{\begin{array}[]{lll}t_{n}+c_{n}t_{\ell},&\quad\text{if}~{}1\leq n\leq\ell-1,\\ t_{\ell},&\quad\text{if}~{}n=\ell.\end{array}\right.$
@@
$\frac{\partial}{\partial t_{n}}~{}\rightarrow~{}\left\{\begin{array}[]{lll}\frac{\partial}{\partial t_{n}},&\quad\text{if}~{}1\leq n\leq\ell-1,\\ \frac{\partial}{\partial t_{\ell}}+\mathop{\textstyle\sum}\limits_{i=1}^{\ell-1}c_{i}\frac{\partial}{\partial t_{i}}&\quad\text{if}~{}n=\ell.\end{array}\right.$
@@ 3.2. Non-crossing permutations and the main theorem
Here we consider several values of $\alpha\in\mathbb{R}$, say $\alpha_{k}$ for $k=1,\ldots,K$. For each $\alpha=\alpha_{k}$, let $\mathcal{I}[\alpha_{k}]$ be a subset of []=$[\ell]=\{1,\ldots,\ell\}$ with $|\mathcal{I}[\alpha_{k}]|=m_{k}\leq\ell$. We then consider $\Phi_{\ell}(\kappa,\alpha_{k})=0$ in (3.1) for $k=1,\ldots,K$, and define the soliton parameter $\kappa\in\mathbb{R}^{M}$, which is the ordered set of the roots of $\Phi_{\ell}(\kappa,\alpha_{k})=0$,
@@
(3.5) ^$\hat{\mathcal{I}}[\alpha_{k}]:=\phi(\mathcal{I}[\alpha_{k}]):=\{\,m=\phi(j,k):j\in\mathcal{I}[\alpha_{k}]\,\}.$
@@
where $N_{k}=n_{1}+\cdots+n_{k}$ with $11\leq n_{k}\leq m_{k}-$ 1. Note that $E_{j}(\mathbf{t})=\exp(\mathop{\textstyle\sum}\limits_{n=1}^{\infty}\kappa_{j}^{n}t_{n})$. Here we also take $A[\alpha_{k}]:=(a_{i,j}(\alpha_{k}))\in\text{Gr}(n_{k},m_{k})_{\geq 0}$, and let $\pi(A[\alpha_{k}])$ be the corresponding permutation in the cycle notation,
@@ A $\kappa$-direct sum of the matrices $A[\alpha_{k}]$ for $k=1,\ldots,K$ is defined by
^$\widehat{\bigoplus}_{k=1}^{K}A[\alpha_{k}]:=\left(a_{i,m}\right)_{1\leq i\leq N,~{}1\leq m\leq M},$
@@
^$\hat{\mathcal{I}}[\alpha_{1}]=\{1,5,6,9\},\qquad\hat{\mathcal{I}}[\alpha_{2}]=\{2,3,4,7,8\}.$
@@
    Two cycles $\pi^{(p)}$ and $\pi^{(q)}$ are non-crossing, if the corresponding chord diagrams have no crossing chords between two permutations.
@@
    Two matrices $A[\alpha_{i}]$ and $A[\alpha_{j}]$ for $i\neq j$ are non-crossing, if the corresponding permutations $\pi(A[\alpha_{i}])$ and $\pi(A[\alpha_{j}])$ are noncrossing.
@@
Proof. Since $A[\alpha_{i}]$ and $A[\alpha_{j}]$ are non-crossing, we have from Lemma 3.5 that the number of free parameters of $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$ is just the sum of those in $A[\alpha_{i}]$ and $A[\alpha_{j}]$, and the number of free parameters in the totally nonnegative matrix $A[\alpha_{i},\alpha_{j}]$ is the same as that of $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$. These free parameters are given by the nonzero entires of the matrices $A[\alpha_{i}]$ and $A[\alpha_{j}]$. Let $a_{k,l}$ be a nonzero entry corresponding to that in $A[\alpha_{i}]$ or $A[\alpha_{j}]$, where $\{{k},{l}\}$ is either in ^$\hat{\mathcal{I}}[\alpha_{i}]$ or ^$\hat{\mathcal{I}}[\alpha_{j}]$. Then there exists a unique pair $(I,J)\in\mathcal{M}(A[\alpha_{i}])\times\mathcal{M}(A[\alpha_{j}])$, so that $\Delta_{\hat{I}\cup\hat{J}}(A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}])=\pm\Delta_{I}(A[\alpha_{i}])\Delta_{J}(A[\alpha_{j}])=\pm a_{k,l}$, where $\hat{I}=\phi(I)$ and $\hat{J}=\phi(J)$. Note also that $\Delta_{\hat{I}\cup\hat{J}}(A[\alpha_{i},\alpha_{j}])=\pm a_{k,l}$. This determines the signs of all nonzero entries $a_{{k},{l}}$ in $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$. Since $A[\alpha_{i},\alpha_{j}]$ is unique and have the same sets of minors as $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$, we have that $\Delta_{\hat{I}\cup\hat{J}}(A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}])>0$ for all $I\in\mathcal{M}(A[\alpha_{i}])$ and $J\in\mathcal{M}(A[\alpha_{j}])$
@@
Suppose that $A[\alpha_{i}]$ and $A[\alpha_{j}]$ are non-crossing for any $1\leq i\neq j\leq K$. Then we can make the combined matrix ^$\hat{\oplus}_{k=1}^{K}A[\alpha_{k}]$ totally nonnegative, and we have
@@
where $A[\alpha_{1},\ldots,\alpha_{K}]$ denotes the totally nonnegative matrix associated to ^$\hat{\oplus}_{k=1}^{K}A[\alpha_{k}]$, i.e. $A[\alpha_{1},\ldots,\alpha_{K}]\in\text{Gr}(N,M)_{\geq 0}$ with $N=n_{1}+\cdots+n_{K}$ and $M=m_{1}+\cdots+m_{K}$.

### v2

@@ 1. Introduction
The KP equation is integrable, and has infinitely many symmetries represented by the commuting flows. The parameters of the flows are expressed by $\{t_{j}:j=1,2,\ldots\}$, in particular, we denote $t_{1}=x,t_{2}=y,t_{3}=t$ for the KP equation. The set of all the flows defines the KP hierarchy, whose first member is the KP equation, and the variable $u$ is now considered as a function of $\mathbf{t}=(t_{1},t_{2},\ldots)$. The real regular soliton solutions of the KP equation, referred to as the KP solitons, has been recently classified in [5, 12] (see the book [10] for the review of these results). Each KP soliton can be determined by a pair of two real data $(\kappa,A)$, referred to as soliton parameters, where $\kappa=(\kappa_{1},\ldots,\kappa_{M})\in\mathbb{R}^{M}$ and an $N\times M$ matrix $A\in\text{Gr}(N,M)_{\geq 0}$, the totally nonnegative (tnn) Grassmannian defined by the set of $N$-dimensional subspaces in $\mathbb{R}^{M}$ with all nonnegative maximal minors of $A$. We also recall that the element $A\in\text{Gr}(N,M)_{\geq 0}$ can be uniquely parameterized by the derangement (permutation without fixed point, i.e. no 1-cycle) of symmetric group $S_{M}$, denoted by $\pi(A)\in S_{M}$. Then the soliton solution $u(x,y,t)$ is given in terms of the $\tau$-function (see e.g. [10]),
@@
where 2-cycle $(i,j)$ implies the transposition $i\leftrightarrow j$. Note here that all the 2-cycles in (1.4) form a nesting (see below, also [6] for the definition). Each soliton with the parameter $\kappa=(\kappa_{k},-\kappa_{k})$ in the $N$-soliton solution is represented by the transposition (k,2Nk+1)(k,2N-k+1), which is considered as the derangement associated to $A_{k}\in\text{Gr}(1,2)_{\geq 0}$, i.e. $\pi(A_{k})=(k,2N-k+1)$ and $\pi(A)=\prod_{k}\pi(A_{k})$.
@@
which is the spectral curve of solitons for the 3-reduction. As one of the main results (Theorem 4.2), we give the regular soliton solutions of the Boussinesq equation in (1.7), whose soliton matrix $A$ is parametrized by the derangement in the cycle notation either
@@
where each $\kappa_{j}[\alpha_{i}]$ is a root of $\Phi_{\ell}(\kappa,\alpha_{i})=0$, and $\mathcal{I}[\alpha_{i}]$ is a subset of $\ell]:=\{1,\ldots,\ell\}$ with $|\mathcal{I}[\alpha_{i}]|=m_{i}$. Here $M=m_{1}+\cdots+m_{K}$. For each $m_{i}$, take $1\leq n_{i}\leq m_{i}-1$, and set $N=n_{1}+\cdots+n_{K}$. Then for each pair $(n_{i},m_{i})$, construct the matrix element $A[\alpha_{i}]\in\text{Gr}(n_{i},m_{i})_{\geq 0}$, so that the corresponding derangements $\pi(A[\alpha_{i}])$ are mutually non-crossing for $i=1,\ldots,K$. Then the regular soliton solution of the $\ell$-reduction is given by the soliton parameters $\kappa\in\mathbb{R}^{M}$ and $A[\alpha_{1},\ldots,\alpha_{K}]\in\text{Gr}(N,M)_{\geq 0}$ associated with a ($\kappa$-)direct sum of the matrices $A[\alpha_{i}]$,
${\widehat{\bigoplus}}_{i=1}^{K}A[\alpha_{i}],$
@@ 2.1. The KP hierarchy
where $p_{i}({\bf x})$ is the elementary Schur polynomial of degree $i$ and $\tilde{\partial}=(\partial_{1},\frac{1}{2}\partial_{2},\frac{1}{3}\partial_{3},\ldots)$, which is defined by
@@ 2.3. KP solitons
where $\kappa\in\mathbb{R}^{M}$ whose elements are ordered as $(\kappa_{1}<\cdots<\kappa_{M})$, and $A=(a_{i,j})$ is an $N\times M$ real matrix with full rank. Let $I=\{i_{1}<i_{2}<\cdots<i_{N}\}$ be an $N$ ordered subset of the index set $M]:=\{1,2,\ldots,M\}$. Then the $\tau$-function (2.6) becomes (1.3), and using the Binet-Cauchy lemma, the $\tau$-function can be expressed by
@@
    For $y\gg 0$, there are $i_{n},\pi(i_{n})]-$solitons with $\pi(i_{n})>i_{n}$ for $n=1,\ldots,N$.
@@
    For $y\ll 0$, there are $\pi(j_{m}),j_{m}]-$solitons with $\pi(j_{m})<j_{m}$ for $m=1,\ldots,M-N$.
@@ 2.4. The Gel’fand-Dickey $\ell$-reductions
that is, all the variables $\{r_{i}:1\leq i\leq\ell-1\}$ do not depend on the times $t_{n\ell}$. Following [14], we define the $\ell$-reduction as follows. Let $\{n_{1},\ldots,n_{K}\}$ be a partition of $N$ with $1\leq n_{k}<\ell$, i.e. $N=n_{1}+\cdots+n_{K}$, and define $N_{k}=n_{1}+\cdots+n_{k}$ $(\text{i.e.}~N=N_{K})$. Then the $\ell$-reduction is defined by a condition for the functions $\{f_{i}:1\leq i\leq N\}$ in (2.7) given by
@@
One should note that the reduction equation (2.12) implies that we have $f_{i}(\mathbf{t})=e^{\mathop{\textstyle\sum}\limits_{n}\alpha_{k}^{n}t_{n\ell}}g_{i}(\mathbf{t}^{\prime})$ with $\mathbf{t}^{\prime}=(t_{j}:j\neq 0~(\text{mod}~\ell))$, which shows that the $\tau$-function is given by $\tau(\mathbf{t})=\text{Wr}(f_{1},\ldots f_{N})(\mathbf{t})=e^{\mathop{\textstyle\sum}\limits_{k}\mathop{\textstyle\sum}\limits_{n}\alpha_{k}^{n}t_{n\ell}}\text{Wr}(g_{1},\ldots,g_{N})(\mathbf{t}^{\prime})$. Then the variables $r_{n}(\mathbf{t})$ in (2.11) has no dependency on the flow parameters $t_{n\ell}$.
@@ 3.1. The spectral curve and the $\tau$-function
(3.2) $t_{n}~\rightarrow~\left\{\begin{array}[]{lll}t_{n}+c_{m}t_{k\ell},&\quad\text{if}~n=m+(k-1)\ell\quad\text{with}\quad 1\leq m\leq\ell-1\quad\text{and}\quad k\geq 1,\\ t_{k\ell},&\quad\text{if}~n=k\ell\quad\text{with}\quad k\geq 1.\end{array}\right$.
@@
$\frac{\partial}{\partial t_{n}}~\rightarrow~\left\{\begin{array}[]{lll}\frac{\partial}{\partial t_{n}},&\quad\text{if}~n\neq k\ell\quad\text{with}\quad k\geq 1,\\ \frac{\partial}{\partial t_{k\ell}}+\mathop{\textstyle\sum}\limits_{m=1}^{\ell-1}c_{m}\frac{\partial}{\partial t_{(k-1)\ell+m}},&\quad\text{if}~n=k\ell\quad\text{with}\quad k\geq 1.\end{array}\right$.
@@ 3.2. Non-crossing permutations and the main theorem
Here we consider several values of $\alpha\in\mathbb{R}$, say $\alpha_{k}$ for $k=1,\ldots,K$. For each $\alpha=\alpha_{k}$, let $\mathcal{I}[\alpha_{k}]$ be a subset of $\ell]=\{1,\ldots,\ell\}$ with $2\leq|\mathcal{I}[\alpha_{k}]|=m_{k}\leq\ell$. We then consider $\Phi_{\ell}(\kappa,\alpha_{k})=0$ in (3.1) for $k=1,\ldots,K$, and define the soliton parameter $\kappa\in\mathbb{R}^{M}$, which is the ordered set of the roots of $\Phi_{\ell}(\kappa,\alpha_{k})=0$,
@@
(3.5) $\hat{\mathcal{I}}[\alpha_{k}]:=\phi(\mathcal{I}[\alpha_{k}]):=\{\,m=\phi(j,k):j\in\mathcal{I}[\alpha_{k}]\,\}.$
@@
where $N_{k}=n_{1}+\cdots+n_{k}$ with $1\leq n_{k}\leq m_{k}-1$. Note that $E_{j}(\mathbf{t})=\exp(\mathop{\textstyle\sum}\limits_{n=1}^{\infty}\kappa_{j}^{n}t_{n})$. Here we also take $A[\alpha_{k}]:=(a_{i,j}(\alpha_{k}))\in\text{Gr}(n_{k},m_{k})_{\geq 0}$, and let $\pi(A[\alpha_{k}])$ be the corresponding permutation in the cycle notation,
@@ A $\kappa$-direct sum of the matrices $A[\alpha_{k}]$ for $k=1,\ldots,K$ is defined by
$\widehat{\bigoplus}_{k=1}^{K}A[\alpha_{k}]:=\left(a_{i,m}\right)_{1\leq i\leq N,~1\leq m\leq M},$
@@
$\hat{\mathcal{I}}[\alpha_{1}]=\{1,5,6,9\},\qquad\hat{\mathcal{I}}[\alpha_{2}]=\{2,3,4,7,8\}.$
@@
    Two permutations $\pi^{(p)}$ and $\pi^{(q)}$ are non-crossing, if the corresponding chord diagrams have no crossing chords between two permutations.
@@
    Two matrices $A[\alpha_{i}]$ and $A[\alpha_{j}]$ for $i\neq j$ are non-crossing, if the corresponding permutations $\pi(A[\alpha_{i}])$ and $\pi(A[\alpha_{j}])$ are non-crossing.
@@
Proof. Since $A[\alpha_{i}]$ and $A[\alpha_{j}]$ are non-crossing, we have from Lemma 3.5 that the number of free parameters of $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$ is just the sum of those in $A[\alpha_{i}]$ and $A[\alpha_{j}]$, and the number of free parameters in the totally nonnegative matrix $A[\alpha_{i},\alpha_{j}]$ is the same as that of $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$. These free parameters are given by the nonzero entires of the matrices $A[\alpha_{i}]$ and $A[\alpha_{j}]$. Let $a_{k,l}$ be a nonzero entry corresponding to that in $A[\alpha_{i}]$ or $A[\alpha_{j}]$, where $\{{k},{l}\}$ is either in $\hat{\mathcal{I}}[\alpha_{i}]$ or $\hat{\mathcal{I}}[\alpha_{j}]$. Then there exists a unique pair $(I,J)\in\mathcal{M}(A[\alpha_{i}])\times\mathcal{M}(A[\alpha_{j}])$, so that $\Delta_{\hat{I}\cup\hat{J}}(A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}])=\pm\Delta_{I}(A[\alpha_{i}])\Delta_{J}(A[\alpha_{j}])=\pm a_{k,l}$, where $\hat{I}=\phi(I)$ and $\hat{J}=\phi(J)$. Note also that $\Delta_{\hat{I}\cup\hat{J}}(A[\alpha_{i},\alpha_{j}])=\pm a_{k,l}$. This determines the signs of all nonzero entries $a_{{k},{l}}$ in $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$. Since $A[\alpha_{i},\alpha_{j}]$ is unique and have the same sets of minors as $A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}]$, we have that $\Delta_{\hat{I}\cup\hat{J}}(A[\alpha_{i}]\hat{\oplus}A[\alpha_{j}])>0$ for all $I\in\mathcal{M}(A[\alpha_{i}])$ and $J\in\mathcal{M}(A[\alpha_{j}])$.
@@
Suppose that $A[\alpha_{i}]$ and $A[\alpha_{j}]$ are non-crossing for any $1\leq i\neq j\leq K$. Then we can make the combined matrix $\hat{\oplus}_{k=1}^{K}A[\alpha_{k}]$ totally nonnegative, and we have
@@
where $A[\alpha_{1},\ldots,\alpha_{K}]$ denotes the totally nonnegative matrix associated to $\hat{\oplus}_{k=1}^{K}A[\alpha_{k}]$, i.e. $A[\alpha_{1},\ldots,\alpha_{K}]\in\text{Gr}(N,M)_{\geq 0}$ with $N=n_{1}+\cdots+n_{K}$ and $M=m_{1}+\cdots+m_{K}$.
@@
Remark 3.9.
In the combinatorics of totally nonnegative Grassmannian, non-crossing permutations play a fundamental role in totally nonnegativity, as explained in [1]. It may be interesting to establish a precise connection between our results and their results.